OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P? (x – 2)2 + (y + 1)2 = 64 (x + 2)2 + (y – 1)2 = 64 (x – 2)2 + (y + 1)2 = 8 (x + 2)2 + (y – 1)2 = 8

OpenStudy (callisto):

centre - radius form : (x – h)^2 + (y + k)^2 = r^2 where (h,k) is the centre of circle r is the radius can you try it now?

OpenStudy (anonymous):

(x-2)^2 + (y+-1)^2=r^2 4+2=12?

OpenStudy (callisto):

sorry centre - radius form : (x – h)^2 + (y - k)^2 = r^2

OpenStudy (callisto):

(x-2)^2 <- correct

OpenStudy (callisto):

r = radius , that is given in the question..

OpenStudy (callisto):

can you try again using the right one ?

OpenStudy (anonymous):

4+(y-k)^2=r^2 4+2=6 = 16 =96?

OpenStudy (callisto):

nope...

OpenStudy (callisto):

you just need to put h = 2, k =-1 and r=8 into the equation (x – h)^2 + (y - k)^2 = r^2 can you try again?

OpenStudy (anonymous):

(x-2)+(y--1)=8^2= Am I supposed to multiply them by 2 because that's what I'm doing and it keeps coming out wrong.

OpenStudy (callisto):

hmm.. 8^2 = 8*8 = 64

OpenStudy (callisto):

\[8^2 \neq 8\times2\]